OpenStudy (anonymous):

Use identities (no calculators) to fi nd the exact value for (sin 9)(sin 36)-(cos 9)(cos 36)

OpenStudy (anonymous):

9+36 =?

OpenStudy (anonymous):

how can you play with those angles to get an angle for which there is an exact value

OpenStudy (anonymous):

does this equals cos 45?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is 1/ sqrt(2)

OpenStudy (anonymous):

\[\sin \theta = \cos 90 - \theta\]

OpenStudy (anonymous):

How do I figure out if its positive of negative

OpenStudy (anonymous):

sines, cosines and tangents in the first quadrant (0-90 degrees) are all positive

OpenStudy (anonymous):

i've just checked my maths formula book and the given expression = -cos(9 + 36) not cos (9+36) so the correct answer is -(1/sqrt2)

OpenStudy (anonymous):

this is not the formula for \[cos(a+b)\] but it is its negative.

OpenStudy (anonymous):

of course 9+36=36+9=45

OpenStudy (anonymous):

but the formula for \[cos(a+b)=cos(a)cos(b)-sin(a)sin(b)\]

OpenStudy (anonymous):

and \[sin(a)sin(b)-cos(a)cos(b)=-(cos(a)cos(b)-sin(a)sin(b)\] that is why you had to change the sign from \[\frac{\sqrt{2}}{2}\]to \[-\frac{\sqrt{2}}{2}\]